How many liters of 20% salt solution must be added to 60 liters of 40% salt solution to get a solution that is 35% salt?
Answer by Earlsdon(6294)

Let x = the number of liters of 20% salt solution to be added to 60 liters of 40% salt solution to obtain (60+x) liters of 35% salt solution.
You can write the equation to solve for x, after converting the percentages to their decimal equivalents:
0.2x+0.4(60) = 0.35(60+x) Simplify.
0.2x+24 = 21+0.35x Subtract 0.2x from both sides.
24 = 21+0.15x Subtract 21 from both sides.
3 = 0.15x Divide both sides by 0.15
20 = x
You would need to add 20 liters of 20% salt solution to 60 liters of salt solution to obtain 80 liters of 35% salt solution.